22 out of 58 channels had commercials on
Find the probability of getting a value of the statistic that is smaller than or equal to 22.
(pˆ

Answer by stanbon(75887)

22 out of 58 channels had commercials on
Find the probability of getting a value of the statistic that is smaller than or equal to 22.
----
sample proportion = 22/58 = 0.38
----
margin of error: sqrt[0.38*0.62/58] = 0.064
-----
Statistic <= 22 ??
That doesn't make any sense.
================================
Cheers,
Stan H.